# Exact black hole solutions in gravity with a background Kalb-Ramond field

### v1

@@ I INTRODUCTION
Several theories of quantum gravity have been proposed Birrell and Davies (1984); Maldacena (1998); Aharony et al. (2000); Gubser et al. (1998); Alfaro et al. (2000, 2002); Rovelli and Smolin (1990), but lots of quantum gravity effects can be observed only at the Planck scale ($\sim 10^{19}$ GeV), which is beyond the reach of current experimental capabilities. Some of these quantum gravity theories assume that Lorentz symmetry might be broken in the gravitational UV regime which is at our presently accessible low-energy scales.
@@
In the context of Bumblebee gravity, Casana et al. obtained an exact solution for a static spherically symmetric spacetime Casana et al. (2018). Subsequently, Maluf et al. derived an (A)dS-Schwarzschild-like solution by relaxing the vacuum conditions Maluf and Neves (2021), while Xu et al. identified new classes of static spherically symmetric Bumblebee black holes with a nonzero temporal component of the bumblebee vector field and investigated their thermodynamic properties as well as observational implications Xu et al. (2023). Ding et al. obtained a slowly rotating Bumblebee black hole solution and a BTZ-like black hole solution Ding et al. (2020, 2023). Moreover, several new black hole solutions were recently proposed within the framework of Bumblebee gravity Liu et al. (2025a); Ding et al. (2024). Different properties of these black holes were extensively explored in various works Kuang and Övgün (2022); Oliveira et al. (2019); Liu et al. (2023); Gomes et al. (2020); Kanzi and Sakallı (2019); Güllü and Övgün (2022); Oliveira et al. (2021); Kanzi and Sakallı (2021); Ovgün et al. (2018); Sakallı and Yörük (2023); Mangut et al. (2023).
@@
In another Lorentz violation model, the Kalb-Ramond field, which arises from the bosonic spectrum of string theory Kalb and Ramond (1974), is a self-interacting second-rank antisymmetric tensor field. When the tensor field’s nonzero VEV is nonminimally coupled with the gravity sector, it leads to spontaneous Lorentz symmetry violation. This coupling results in the ground state of a physical quantum system being characterized by nontrivial VEVs. Similarly, several black hole solutions have also been proposed and studied in the context of the Kalb-Ramond field Do and Kao (2020); Liu et al. (2024a); Duan et al. (2024); Yang et al. (2023); Lessa et al. (2020); Kumar et al. (2020); Liu et al. (2024b). The corresponding properties of the black holes were also studied Du et al. (2025); Guo et al. (2024); Junior et al. (2024); Jumaniyozov et al. (2024); Ma et al. (2024); al Badawi et al. (2024); Zahid et al. (2024a); Wei et al. (2015); Zahid et al. (2024b); Atamurotov et al. (2022); Lessa et al. (2021); Maluf and Muniz (2022); Liu et al. (2025b); Jha (2024); Fathi and Övgün (2025).
@@
It is worth noting that some previous studies have treated the variation of the nonminimal coupling term between the Kalb–Ramond field and the scalar curvature, $\xi_{1}B^{\mu\nu}B_{\mu\nu}R$, as effectively transforming into $\mp\xi_{1}b^{2}R$ in the vacuum. As a result, they argued that this term could be absorbed into the action via a redefinition of the gravitational coupling constant, making it appear as if the term vanishes from the action. However, this interpretation is misleading. In fact, the variation of $\xi_{1}B^{\mu\nu}B_{\mu\nu}R$ with respect to the metric is not equivalent to the variation of $\mp\xi_{1}b^{2}R$. Therefore, the contribution of this term to the gravitational field equations has been overlooked in such treatments.
@@ II THE KALB-RAMOND MODEL FOR SPONTANEOUS LORENTZ SYMMETRY BREAKING
$S$ $=$ $\int d^{D}x\sqrt{-g}\bigg{[}\frac{1}{2\kappa}(R-2\Lambda)-\frac{1}{12}H^{\mu\nu\rho}H_{\mu\nu\rho}-V(B^{\mu\nu}B_{\mu\nu}\pm b^{2})+\frac{1}{2\kappa}(\xi_{1}B^{\mu\nu}B_{\mu\nu}R$ (1)
@@
with a nonzero VEV $\langle B_{\mu\nu}\rangle=b_{\mu\nu}$. Consequently, it follows that the VEV is constrained by the condition:
@@
Consequently, the VEV of the Kalb-Ramond field leads to two independent background vectors, in contrast to the single vector generated by the VEV of the bumblebee field. In this work, we assume that the only nonvanishing component of the Kalb–Ramond field is the pseudo-electric field, which is given by
$b_{2}=-\tilde{E}(r)dt\wedge dr.$ (8)
Equivalently, it can be expressed in component form as
$b_{10}=-b_{01}=\tilde{E}(r).$ (9)
@@
-- $2B^{\alpha}_{\leavevmode\nobreak\ \mu}B_{\nu}{}_{\alpha}R-\nabla_{\mu}\nabla_{\nu}(B^{\alpha\beta}B_{\alpha\beta})\big{]}\sqrt{-g}.$
@@
$\frac{\delta(b^{2}R\sqrt{-g})}{\delta g^{\mu\nu}}=\xi_{1}b^{2}G_{\mu\nu}\sqrt{-g}.$ (11)
@@
$+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha\beta}B_{\nu\beta}\right)+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha\beta}B_{\mu\beta}\right)-\frac{1}{2}\nabla^{\alpha}\nabla_{\alpha}\left(B_{\mu}^{\gamma}B_{\nu\gamma}\right)-\frac{1}{2}g_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\left(B^{\alpha\gamma}B^{\beta}{}_{\gamma}\right)\bigg{]}$
@@ III STATIC SPHERICALLY SYMMETRIC BLACK HOLE SOLUTIONS
${d}{s}^{2}=-A(r){dt}^{2}+S(r){dr}^{2}+r^{2}{\leavevmode\nobreak\ d}\Omega^{2},$ (14)
where $d\Omega^{2}={\leavevmode\nobreak\ d}\theta^{2}+\sin^{2}\theta{d}\varphi^{2}$.
@@ III.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
$=$ $\frac{48M^{2}}{r^{6}}-\frac{8\ell_{2}M}{(1+\ell_{1}-\frac{\ell_{2}}{2})r^{5}}+\frac{{\ell_{2}}^{2}M^{2}}{(1+\ell_{1}-\frac{\ell_{2}}{2})^{2}r^{4}}.$
This indicates that Lorentz-violating effects are intrinsic to the theory and cannot be removed by any coordinate transformation. The explicit breaking of Lorentz symmetry sets this framework apart from many other modified gravity theories. It is evident that the metric functions approach to $A(r)\rightarrow\frac{1}{1-\gamma}$ and $S(r)\rightarrow 1-\gamma$ at infinity, indicating that the spacetime is not asymptotically Minkowski.
@@ III.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
$V(X)=\frac{\lambda}{2}X,$ (24)
where $\lambda$ is a Lagrange multiplier field Bluhm et al. (2008). Note that the equation of motion for the Lagrange multiplier ensures the vacuum condition $X=0$, which also ensures that $V=0$ for any on-shell field $\lambda$. However, unlike the vacuum condition in Case A, $V^{\prime}\neq 0$ is valid when the field $\lambda$ is nonzero. Thus, the potential $V$ contributes to the field equations. The choice $V(X)=\frac{\lambda}{2}X$ effectively freezes motion about the potential minimum and hence permits an efficient extraction of the essential physics Kostelecky and Samuel (1989b).
Combining the spherically symmetric metric and the form of the potential $V$, we can derive the specific field equations. An analytical solution is possible if and only if
$\lambda=\frac{(4\xi_{1}+\xi_{2})\Lambda}{\kappa(1-\ell_{1}-\frac{\ell_{2}}{2})}.$ (25)
Substituting the above conditions given in Eqs. (24) and (25) back into the field Eqs. (16) - (18), we can obtain
@@
It is easy to see that this solution is similar to the Schwarzschild-(A)dS-like black solution obtained by Yang et al. Yang et al. (2023). From this, we can see that when considering the nonminimal coupling of the Kalb-Ramond field with both the Ricci tensor and the scalar curvature, both contributions affect the resulting black hole solutions.
@@ IV ROTATING BTZ-LIKE BLACK HOLE SOLUTION
$g_{tt}=-A(r)+r^{4}K(r)^{2}=M+\frac{\Lambda}{(1+{\ell})}r^{2}+\frac{\ell j^{2}}{4(1+\ell)r^{2}}.$ (32)
@@
$r_{\pm}=\frac{1}{2}\sqrt{\frac{1}{-\Lambda}}\left(\sqrt{(1+{\ell})M+\sqrt{-\Lambda J}}\pm\sqrt{(1+{\ell})M-\sqrt{-\Lambda J}}\right).$ (34)
@@ V THERMODYNAMICS
where $\bm{\epsilon}$ denotes the spacetime volume form and $L$ corresponds to the action (1). Varying the fields $\Phi\equiv\{g_{ab},B_{ab}\}$ gives
@@
Furthermore, by the Poincaré lemma, this implies that at least locally there exists a $(D-2)$-form, called the Noether charge $\mathbf{Q}_{\xi}$, such that
@@ V.1 Three-Dimensional Case ($D=3$)
${E_{R}}^{abcd}$ $=\frac{1}{2\kappa}\left(X^{abcd}+\xi_{2}B^{\mu\rho}{B^{\nu}}_{\rho}{Y_{\mu\nu}}^{abcd}+\xi_{3}B^{\mu\nu}B_{\mu\nu}X^{abcd}\right),$ (52)
@@
$T_{H}=\frac{\kappa}{2\pi}=\frac{r_{h}^{4}\Lambda_{e}-16j^{2}}{2\pi r_{h}^{3}(1+\ell)}.$ (57)
using the Killing vector $\partial_{t}$ and $\partial_{\phi}$, we obtain from Eq. (45) the expressions for the energy and angular momentum of the black hole:

### v2

@@ I INTRODUCTION
Several theories of quantum gravity have been proposed Birrell and Davies (1984); Maldacena (1998); Aharony et al. (2000); Gubser et al. (1998); Alfaro et al. (2000, 2002); Rovelli and Smolin (1990), but lots of quantum gravity effects can be observed only at the Planck scale ($\sim 10^{19}$ GeV), which is beyond the reach of current experimental capabilities. Some of these quantum gravity theories assume that Lorentz symmetry might be broken in the gravitational IR regime, which corresponds to our presently accessible low-energy scales.
@@
In the context of Bumblebee gravity, Casana et al. obtained an exact solution for a static spherically symmetric spacetime Casana et al. (2018). Subsequently, Maluf et al. derived an (A)dS-Schwarzschild-like solution by relaxing the vacuum conditions Maluf and Neves (2021), while Xu et al. identified new classes of static spherically symmetric Bumblebee black holes with a nonzero temporal component of the bumblebee vector field and investigated their thermodynamic properties as well as observational implications Xu et al. (2023). Ding et al. obtained a slowly rotating Bumblebee black hole solution and a BTZ-like black hole solution Ding et al. (2020, 2023). Moreover, several new black hole solutions were recently proposed within the framework of bumblebee gravity Liu et al. (2025a); Ding et al. (2024). Different properties of these black holes were extensively explored in various works Filho et al. (2023); Kuang and Övgün (2022); Oliveira et al. (2019); Liu et al. (2023); Gomes et al. (2020); Kanzi and Sakallı (2019); Güllü and Övgün (2022); Oliveira et al. (2021); Kanzi and Sakallı (2021); Ovgün et al. (2018); Sakallı and Yörük (2023); Mangut et al. (2023); Araújo Filho et al. (2024).
@@
In another Lorentz violation model, the Kalb-Ramond field $B_{\mu\nu}$, which arises from the bosonic spectrum of string theory Kalb and Ramond (1974), is a self-interacting second-rank antisymmetric tensor field. When the tensor field’s nonzero VEV is nonminimally coupled with the gravity sector, it leads to spontaneous Lorentz symmetry violation. This coupling results in the ground state of a physical quantum system being characterized by nontrivial VEVs. Similarly, several black hole solutions have also been proposed and studied in the context of the Kalb-Ramond field Do and Kao (2020); Liu et al. (2024a); Duan et al. (2024); Yang et al. (2023); Lessa et al. (2020); Kumar et al. (2020); Liu et al. (2024b). The corresponding properties of the black holes were also studied Du et al. (2025); Guo et al. (2024); Junior et al. (2024); Jumaniyozov et al. (2024); Ma et al. (2024); al Badawi et al. (2024); Zahid et al. (2024a); Wei et al. (2015); Zahid et al. (2024b); Atamurotov et al. (2022); Lessa et al. (2021); Maluf and Muniz (2022); Liu et al. (2025b); Jha (2024); Fathi and Övgün (2025); Ghosh et al. (2023); Filho et al. (2024); Araújo Filho (2025); Nair et al. (2024).
@@
It is worth noting that some previous studies have treated the variation of the nonminimal coupling term between the Kalb–Ramond field and the scalar curvature, $\xi_{1}B^{\mu\nu}B_{\mu\nu}R$, as effectively transforming into $\mp\xi_{1}b^{2}R$ in the vacuum. Here, $b_{\mu\nu}$ denotes the vacuum expectation value of $B_{\mu\nu}$, and $b^{\mu\nu}b_{\mu\nu}=\mp b^{2}$. As a result, they argued that this term could be absorbed into the action via a redefinition of the gravitational coupling constant, making it appear as if the term vanishes from the action. However, this interpretation is misleading. In fact, the variation of $\xi_{1}B^{\mu\nu}B_{\mu\nu}R$ with respect to the metric is not equivalent to the variation of $\mp\xi_{1}b^{2}R$. Therefore, the contribution of this term to the gravitational field equations has been overlooked in such treatments. As will be discussed in more detail around Eq. (9) –(10), this nonminimal coupling plays a crucial role in our analysis.
@@ II THE KALB-RAMOND MODEL FOR SPONTANEOUS LORENTZ SYMMETRY BREAKING
$S$ $=$ $\int d^{D}x\sqrt{-g}\bigg[\frac{1}{2\kappa}(R-2\Lambda)-\frac{1}{12}H^{\mu\nu\rho}H_{\mu\nu\rho}-V(B^{\mu\nu}B_{\mu\nu})+\frac{1}{2\kappa}(\xi_{1}B^{\mu\nu}B_{\mu\nu}R$ (1)
@@
with a nonzero VEV $\langle B_{\mu\nu}\rangle=b_{\mu\nu}$. It follows that the VEV of the Kalb-Ramond field is determined by the condition $V(B^{\mu\nu}B_{\mu\nu}\pm b^{2})=0$, which implies the constraint
@@
Consequently, the VEV of the Kalb-Ramond field leads to two independent background vectors, in contrast to the single vector generated by the VEV of the bumblebee field. In spherical coordinates $(t,r,\theta,\phi)$, we assume that the vacuum configuration of the Kalb–Ramond field has only the following nonvanishing components:
$b_{rt}=-b_{tr}=\tilde{E}(r).$ (8)
@@
- $2B^{\alpha}_{\penalty 10000\ \mu}B_{\nu}{}_{\alpha}R-\nabla_{\mu}\nabla_{\nu}(B^{\alpha\beta}B_{\alpha\beta})\big]\sqrt{-g}.$
@@
$\frac{\delta(b^{2}R\sqrt{-g})}{\delta g^{\mu\nu}}=b^{2}G_{\mu\nu}\sqrt{-g}.$ (10)
@@
[…]
@@ III STATIC SPHERICALLY SYMMETRIC BLACK HOLE SOLUTIONS
${d}{s}^{2}=-A(r){dt}^{2}+S(r){dr}^{2}+r^{2}{\penalty 10000\ d}\Omega^{2},$ (13)
where $d\Omega^{2}={\penalty 10000\ d}\theta^{2}+\sin^{2}\theta{d}\varphi^{2}$.
@@ III.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
It is worth noting that, since $b$ is a constant, the pseudo-electric field $\tilde{E}(r)$ also remains constant. In principle, Lorentz-violating parameter can take negative values, but it is constrained to be a small quantity Yang et al. (2023). Current bounds are
$-1.1\times 10^{-10}\leq\ell\leq 5.4\times 10^{-10}.$ (20)
@@
$=$ $\frac{48M^{2}}{r^{6}}-\frac{8\ell_{2}M}{(1+\ell_{1}-\frac{\ell_{2}}{2})r^{5}}+\frac{{\ell_{2}}^{2}}{(1+\ell_{1}-\frac{\ell_{2}}{2})^{2}r^{4}}.$
This indicates that Lorentz-violating effects are intrinsic to the theory and cannot be removed by any coordinate transformation. The explicit breaking of Lorentz symmetry sets this framework apart from many other modified gravity theories. It is evident that the metric functions approach to $A(r)\rightarrow\frac{1}{1-\gamma}$ and $S(r)\rightarrow 1-\gamma$ at infinity. By performing the coordinate transformations $dt=\sqrt{1-\gamma}\,d\hat{t}$ and $dr=\sqrt{1/(1-\gamma)}\,d\hat{r}$, the asymptotic metric becomes
$ds^{2}=-d\hat{t}^{2}+d\hat{r}^{2}+\frac{1}{1-\gamma}\,\hat{r}^{2}d\Omega^{2}.$ (24)
This shows that the temporal and radial sectors coincide with those of Minkowski spacetime in spherical coordinates, while the angular part acquires a constant factor $1/(1-\gamma)$. And the Ricci scalar evaluates to
$R=-\frac{2\gamma}{(1-\gamma)r^{2}}.$ (25)
Hence, the spacetime is not asymptotically Minkowski. From another perspective, for asymptotically flat spacetimes one usually writes $r=\sqrt{x^{2}+y^{2}+z^{2}}$. As $r\to\infty$ along either timelike or lightlike directions, the metric takes the asymptotic form
$g_{\mu\nu}=\eta_{\mu\nu}+\mathcal{O}(r^{-1}),$ (26)
where $\eta_{\mu\nu}$ denotes the Minkowski metric. However, the asymptotic form of our metric can be written as
$ds^{2}=-d\hat{t}^{2}+(1-\gamma)\,dr^{2}+r^{2}d\Omega^{2},$ (27)
so that the deviation from Minkowski spacetime in the radial component is exactly $\gamma$, i.e., a constant, rather than decaying as $\mathcal{O}(r^{-1})$. This shows that the spacetime is not asymptotically flat.
@@ III.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
$V(\lambda,X)=\frac{\lambda}{2}X,$ (28)
where $\lambda$ is a Lagrange multiplier field Bluhm et al. (2008). The equation of motion obtained by varying with respect to $\lambda$ enforces the vacuum condition $X=0$, which implies $V=0$ for any field $\lambda$ on shell. In this way, the construction effectively freezes motion about the potential minimum and hence permits an efficient extraction of the essential physics Kostelecky and Samuel (1989b). Unlike the vacuum condition in Case A, $V^{\prime}\neq 0$ is valid when the field $\lambda$ is nonzero. Thus, the potential $V$ contributes to the field equations. Combining the spherically symmetric metric and the form of the potential $V$, we can derive the specific field equations. Substituting the conditions given in Eqs. (28) and (29) back into the field equations (15)–(17), an exact analytical solution can be obtained, provided the coupling constants satisfy
$\lambda=\frac{(4\xi_{1}+\xi_{2})\Lambda}{\kappa\left(1-\ell_{1}-\tfrac{\ell_{2}}{2}\right)},$ (29)
which follows directly from the field equations and therefore must hold whenever the latter are satisfied. The exact analytical solution is give
@@
It is easy to see that this solution is similar to the Schwarzschild-(A)dS-like black solution obtained by Yang et al. Yang et al. (2023). From this, we can see that when considering the nonminimal coupling of the Kalb-Ramond field with both the Ricci tensor and the scalar curvature, both contributions affect the resulting black hole solutions. We analyze the asymptotic behavior of the metric, taking into account both signs of the effective cosmological constant. After performing a coordinate transformation analogous to Eq. (24), the metric can be written in the asymptotic form:
$ds^{2}=-(1-\Lambda_{e}\hat{r}^{2})\,d\hat{t}^{2}+\frac{1}{1-\Lambda_{e}\hat{r}^{2}}\,d\hat{r}^{2}+\frac{1}{1-\gamma}\,\hat{r}^{2}d\Omega^{2},$ (31)
where
$\Lambda_{e}=\frac{\Lambda}{(1-\gamma)(1-\ell_{1}-\tfrac{\ell_{2}}{2})}.$ (32)
For $\Lambda_{e}>0$, the spacetime is asymptotically dS, with a cosmological horizon at $\hat{r}\sim 1/\sqrt{\Lambda_{e}}$. For $\Lambda_{e}<0$, the spacetime is asymptotically AdS. The constant factor $1/(1-\gamma)$ in the angular sector corresponds to a uniform rescaling of the boundary sphere. Accordingly, the boundary conformal metric reads
$ds_{\rm bdry}^{2}=-d\hat{\tau}^{2}+\frac{1}{1-\gamma}\,d\Omega^{2}.$ (33)
where $\hat{\tau}$ denotes the time coordinate induced on the conformal boundary. Thus, up to this constant rescaling of the angular coordinates, the spacetime retains its asymptotic (A)dS character.
@@ IV ROTATING BTZ-LIKE BLACK HOLE SOLUTION
$g_{tt}=-A(r)+r^{2}K(r)^{2}=M+\frac{\Lambda}{(1+{\ell})}r^{2}+\frac{\ell j^{2}}{4(1+\ell)r^{2}}.$ (39)
@@
$r_{\pm}=\sqrt{\frac{1}{-2\Lambda}}\left(\sqrt{(1+\ell)M\pm\sqrt{(1+\ell)^{2}M^{2}+\Lambda j^{2}}}\right).$ (41)
In general, an ergoregion appears when $g_{tt}>0$, which requires
$(1+\ell)^{2}M^{2}-\Lambda\ell j^{2}\geq 0.$
Solving $g_{tt}=0$, the ergosurface radii are obtained as
$r_{\rm ergo}^{\pm}=\sqrt{\frac{1}{-2\Lambda}}\left(\sqrt{(1+\ell)M\pm\sqrt{(1+\ell)^{2}M^{2}-\Lambda\ell j^{2}}}\right).$ (42)
In the region $r_{\rm ergo}^{-}<r<r_{\rm ergo}^{+}$ we have $g_{tt}>0$. Since $\ell$ is a small parameter ($|\ell|\ll 1$) and $\Lambda<0$, it follows that $r_{\rm ergo}^{+}>r_{+}$ and $r_{\rm ergo}^{-}<r_{+}$, so the ergoregion extends over $r_{+}<r<r_{\rm ergo}^{+}$. Furthermore, the condition $(1+\ell)^{2}M^{2}-\Lambda\ell j^{2}\geq 0$ is automatically satisfied whenever the horizon exists, i.e., when $(1+\ell)^{2}M^{2}+\Lambda j^{2}\geq 0$.
Therefore, the Lorentz-violating parameter $\ell$ does not remove the ergoregion but modifies its size relative to the standard BTZ case. An ergoregion outside the event horizon $r_{+}$ always persists.
@@ V THERMODYNAMICS
where $\bm{\epsilon}$ denotes the spacetime volume form and $L$ corresponds to the Lagrangian in action (1). Varying the fields $\Phi\equiv\{g_{ab},B_{ab}\}$ gives
@@
Furthermore, by the Poincaré lemma, this implies that at least locally there exists a (D2)(D-2)-form, called the Noether charge $\mathbf{Q}_{\xi}$, such that
@@ V.1 Three-Dimensional Case ($D=3$)
${E_{R}}^{abcd}$ $=\frac{1}{2\kappa}\left(X^{abcd}+\xi_{2}B^{\mu\rho}{B^{\nu}}_{\rho}{Y_{\mu\nu}}^{abcd}+\xi_{1}B^{\mu\nu}B_{\mu\nu}X^{abcd}\right),$ (60)
@@
$T_{H}=\frac{\kappa}{2\pi}=\frac{r_{h}^{4}\Lambda_{e}-16j^{2}}{2\pi r_{h}^{3}(1+\ell)},$ (65)
where $r_{h}$ denotes the horizon radius of the black hole. Using the Killing vector $\partial_{t}$ and $\partial_{\phi}$, we obtain from Eq. (53) the expressions for the energy and angular momentum of the black hole: